Use the defination of inverse function to show analytically that f and g are inverse:
f(x) = 4x+3 and g(x)= x-3/4

Answer by funmath(2873)

Use the defination of inverse function to show analytically that f and g are inverse:
In order to be inverses, their composites have to equal x.
f(x) = 4x+3 and g(x)= (x-3)/4
(fog)(x)=f(g(x))=f((x-3)/4)=4((x-3)/4)+3=x-3+3=x
(gof)(x)=g(f(x))=g(4x+3)=((4x+3)-3)/4=4x/4=x
They are inverses of each other.
Happy Calculating!!!!